OpenStudy (btaylor):

Logarithm questions: how do I switch up the bases? \[2 \log_{4} 9 - \log_{2} 3\]

OpenStudy (anonymous):

\[\log_{a}b=1/(\log_{b}a) \]

OpenStudy (anonymous):

but you dont hv to switch

OpenStudy (anonymous):

\[2\log_2 3-\log_2 3\]

OpenStudy (anonymous):

lets make some sense out of this \[\log_4(x)=\frac{\log_2(x)}{\log_2(4)}\] by the change of base formula and since \(\log_2(4)=2\) we get \[\log_4(x)=\frac{1}{2}\log_2(x)\]

OpenStudy (btaylor):

thanks!